# Parallel Backpropagation for Shared-Feature Visualization

### v1

@@ Abstract
Figure 1: Why does this object image activate IT neurons that are selective for bodies? Our goal is to visually explain reponses of category-selective neurons to outside-of-category (ooc.) stimuli. We start with an ooc. stimulus (object), that strongly activates a neuron which primarily fires for bodies. We compute latent CNN activations for the image, as well as for a set of within-category reference images. A neuron-specific similarity metric operating on the latent activations finds the reference image most similar to the ooc. stimulus. Both images are then passed through the proposed parallel backpropagation method to highlight the shared features driving the neural response.
@@ 1 Introduction
To address this gap, we propose a deep-neural-network-based method for visualizing features of an ooc. stimulus, which are responsible for eliciting high responses from an otherwise category-selective neuron. Our approach relies on analyzing the similarity of the image to a selected within-category image which displays similar features, as shown in Fig. 1. Specifically, after finding a within-category image with similar, neuron-specific features, we use gradient methods to highlight features extracted from a vision model which are shared between the two images and drive the neural response. This helps to visually answer the question why a feature detector that is preferably activated by within-class images, would also respond to the ooc. image in question. The procedure is entirely class-agnostic, and can therefore be used for any selective brain region, as well as for studying internal behaviour of artificial vision models. In this work, we show results for a novel set of multi-unit recordings from body-selective regions in macaque superior temporal sulcus. Body cells constitute a particularly interesting problem, as different body poses produce vast variability among body images, suggesting a rich set of features driving these neurons. We summarize our contributions as follows:
@@
  • We present results from multi-unit recordings from body-selective neurons in macaque IT cortex, demonstrating that these neurons encode overlapping visual features for bodies and objects,
@@ 2 Related work
A substantial body of work exists on attributing behaviour of deep computer vision models to specific image regions. The most common goal is to understand why classification models output the observed class label, given an input image, meaning that attribution methods are often applied to the very last network layer. Several families of approaches have been proposed for tackling this problem: The first relies on image perturbations like occlusion [12, 13], the second is based on an analysis of latent activations [14, 15], and the third computes gradients of class activations w.r.t. pixel intensities [16, 17, 18]. More recent work [19, 20] has shown that some of these methods rely too strongly on the input image itself, being independent of the network to varying degrees. Even though our reweighting scheme is compatible with any attribution method that is computable for latent activations, we therefore rely on vanilla backpropagation in this work, which has been shown to be sensitive to network weights. We provide results for additional attribution methods in the appendix. Further, relevant to our work is that of [21] which computes saliency maps for image similarity. The approach is similar to that of Grad-CAM, as it relies on analyzing feature activations before and after a global pooling layer at the end of the network. For that reason, the latter method is mainly suitable for visualizing global similarity as judged by a network trained to perform this task, rather than more local features relevant for biological neurons along the visual hierarchy in the primate brain.
@@ 3 Methods
3.1 Modelling neural reponses
In recent years, a pretrained CNN backbone combined with a trainable linear readout module has become the gold standard in modelling the stimulus-driven variance in neural responses to images [22, 23]. In these models, an image $x\in\mathbb{R}^{3\times h\times w}$ is fed through a CNN $f(\cdot)$ up to a predetermined layer to obtain a latent representation $a\in\mathbb{R}^{c}$, where $c$ is the dimension of a commonly flattened tensor. The predicted response for a neuron is then computed as
@@ 3.2 Neuron-specific image similarity
If an image $x_{\text{out}}$ strongly drives a neuron, we propose to interpret the relevant image features by studying a visually similar image from the category that the neuron is selective to. To quantify similarity, we adapt the common method of computing the cosine similarity of latent activations computed using a CNN [24]. Since we have additional information on which latent dimensions drive the neuron, we weight feature $i$ by the corresponding weight $w^{(i)}$. Formally, for images $x_{1}$, $x_{2}$ and their corresponding latent activations $a_{1}:=f(x_{1})$, $a_{2}$, we define the neuron-specific image similarity as
@@ 3.3 Parallel backpropagation for shared-feature visualization
In order to further highlight shared features contributing to model activity, we compute weighted gradients of the latent model activations w.r.t. to image pixels [16]. This method was originally for classification models to highlight those image pixels that most strongly influence a given class probability. We build on this work to visualize shared features between images by introducing a simple reweighting procedure. Specifically, we strengthen the influence of features that are shared, while attenuating features that are specific to only one of the images. For the sake of brevity, we illustrate the procedure for the gradients of $x_{\text{out}}$ only, as the process is analogous for $x_{\text{in}}$. First of all, note that the gradient of the predicted neural response $\hat{y}_{\text{out}}$ is given by
@@
Before reweighting the features, we first smooth and normalize each gradient to have unit norm, i.e. $||\frac{\partial}{\partial x_{\text{out}}}a_{\text{out}}^{(i)}||_{2}=1$. Smoothing alleviates pixel-level noise and allows the user to determine regions of pixels with high contribution. It has been shown that a post-processing smoothing step substantially improves the quality of gradient-based attributions [20]. Normalization ensures that the pixel saliency map is not determined solely by the gradient magnitude, as some features may be more sensitive to pixel perturbation than others. More importantly, it allows us to later bound the saliency map from above, which substantially improves interpretability. Subsequently, we weight each feature by an asymmetric variation of its contribution to the neuron-specific similarity metric given in (2). More specifically, we define the weight for feature $i$ by its contribution to $s(\cdot,\cdot)$, given by
@@
The revised weight vector $\beta$ imposes high weights only on those features that are highly activated in the feature vectors of both images (captured by $a_{1}^{(i)}$ and $a_{2}^{(i)}$), and are also relevant for the neuron’s activity (captured by $(w^{(i)})^{2}$). Dividing by the squared norm allows us to bound the intensity of the saliency maps by writing
@@ 4 Experimental setup
For gathering neural data, we utilized two sets of images. The first consisted of 475 images of a macaque avatar on a gray background (see Fig. 1, 2). The images were subsampled from a set of 720 images in which the avatar appeared in 45 unique poses extracted from nine different behavioural classes, each shown from 16 viewpoints [32]. The second set comprised 6,857 objects from varying categories shown on the same gray background and was combined from the OpenImages dataset [33], as well as several smaller ones [34, 35]. To test for body selectivity, we used an additional set of 2068 control body images including a variety of species. These images were only used to test for category-selectivity of the recorded cells. All other references to body images in the text refer to the original image set showing the monkey avatar. All stimuli were centered with respect to the fixation point. They were shown to the monkey at a resolution of 280x280, and were resized to 224 for the Resnet.
@@
We recorded multi-unit activity (MUA) from and surrounding two fMRI-defined body category-selective patches [5] in the macaque superior temporal sulcus (STS), using 16-channel Plexon V probes, while the subjects performed a fixation task. Since body patch neurons typically discharge sparsely for non-body objects, we employed online stimulus selection: In a first phase, we recorded responses to the set of $475$ monkey avatar images. We then trained a model for each channel to predict neural responses to these images. Subsequently, we predicted neural responses to our set of object images, and for each neuron we selected the highest and lowest predicted activator, as well as the object most similar to the top-activating body image, according to $s(\cdot,\cdot)$. We followed the same procedure for the control body images. Finally, in a second experimental phase we recorded responses to these novel images as well as a subset of 75 of the original body images, to test recording stability. We included recording channels in the analysis if the test/retest stability was higher than .60, as measured by the correlation between responses to body images before and after model fitting. Further, we tested each channel for body-category-selectivity by comparing the median response to the selected objects and the selected control bodies using a Mann-Whitney U-test. Further experimental details are given in the appendix.
@@ Conclusion.
We presented a method for visualizing selectivity of class-selective visual feature detectors when confronted with out-of-class images. Further, we showed that body-selective neurons encode bodies and objects using an at least partially shared feature set. We visualized these features, providing an explanation for why some objects activate body-selective neurons. Future work could involve using the same method for other category-selective areas, like face patches. Finally, future work could test these visualizations by presenting highlighted fragments to the subject in a closed-loop fashion.
@@ A.2 Additional details for neural data collection
We recorded neuronal responses during passive fixation task using a 2x2 degree fixation window (EyeLink 1000 infrared eye tracker sampling at 1000 Hz). Stimuli were shown gamma-corrected on a 22.5-inch ViewPixx monitor at a distance of about 57 cm, with a resolution of 1920x1080 and a refresh rate of 120 Hz. A set of 475 stimuli of a monkey avatar in various poses were presented 8 times in a pseudorandom sequence for 200 ms on a gray background, with a 250 ms interstimulus interval. During the interstimulus interval, only a fixation dot was present and monkeys could receive a brief juice reward . Stimulus onset and offset were indicated by a photodiode, detecting luminance changes synced with the stimuli, in the display corner (invisible for the monkeys). Control of stimulus presentation, event timing, and juice delivery was managed by an in-house Digital Signal Processing-based computer system, which also monitored the photodiode signal and tracked eye positions. Neuronal data was collected from and surrounding two fMRI-defined body-selective patches in the ventral superior temporal sulcus (STS) using 16-site linear electrodes (Plexon V probe) with Open Ephys acquisition board and software (sampling rate: 30000 Hz, filtered between 500-5000 Hz). Multi-unit activities were extracted using Plexon Offline Sorter, after applying a high-pass Butterworth filter with a cutoff frequency of 250 Hz. From the stimulus event synchronized continuous data, 550 ms trials were extracted, comprising 200 ms prestimulus and 150 ms poststimulus periods. Responsive (at least for one stimulus showing stronger than 5 spikes/s net responses (baseline: -75 to +25 ms, response: +50 to +250 ms)), and body-selective MUAs (p<0.01 Kruskal-Wallis test), with a split-half reliability >0.5 (Spearman Brown corrected) were selected. The responses of these MUA sites were used to predict neural responses to our set of object and body images, and for each neuron we selected the highest and lowest predicted activator, as well as the object most similar to the top-activating avatar stimuli, according to s(·, ·). Finally, in a second experimental phase we recorded responses to these object and body images as well as a subset of the original monkey avatar stimuli, to test recording stability (same experimental design as in the first phase). For all neurons considered in this work, we tested for body-selectivity by comparing the median response to bodies to the median response to objects using a Mann-Whitney U-test, considering only channels for which the test detected a significant difference.
@@
Two male 7 years old rhesus monkeys (Macaca mulatta) contributed to this study. The animals were housed in enclosures at the KU Leuven Medical School and experienced a natural day-night cycle. Each monkey shared its enclosure with at least one other cage companion. On weekdays, dry food was provided ad libitum, and the monkeys obtained water, or other fluids, during experiments until they were satiated. During weekends, the animals received water along with a mixture of fruits and vegetables. The animals had continuous access to toys and other forms of enrichment. After fMRI scanning, we implanted a custom-made plastic recording chamber, allowing a dorsal approach to temporal body patches. In each animal, the location of the recording chamber was guided by the fMRI body localizer (https://pubmed.ncbi.nlm.nih.gov/36717042/) Surgery was perfomed using standard aseptic procedures and under full anesthesia. Both the animal care and experimental procedures adhere to regional (Flanders) and European guidelines and have been approved by the Animal Ethical Committee of KU Leuven.

### v2

@@ Abstract
Figure 1: Why does this object image activate IT neurons that are selective for bodies? Our goal is to visually explain responses of category-selective neurons to outside-of-category (ooc.) stimuli. We start with an ooc. stimulus (object), that strongly activates a neuron which primarily fires for a specific category (bodies). We compute latent CNN activations for the image, as well as for a set of within-category reference images. A neuron-specific similarity metric operating on the latent activations finds the reference image most similar to the ooc. stimulus. The proposed parallel backpropagation method then highlights the shared features driving the neural response.
@@ 1 Introduction
To address this gap, we propose a deep-neural-network-based method for visualizing features of an ooc. stimulus, which are responsible for eliciting high responses from an otherwise category-selective neuron. Our approach relies on analyzing the similarity of the image to a selected within-category image which displays similar features, as shown in Fig. 1. Specifically, after finding a within-category image with similar, neuron-specific features, we use gradient methods to highlight features extracted from a vision model which are shared between the two images and drive the neural response. This helps to visually answer the question why a feature detector that is preferably activated by within-class images, would also respond to the ooc. image in question. The procedure is compatible with any backbone visualization method that returns a saliency map for hidden units of a convolutional neural network. Further, it is entirely class-agnostic, and can therefore be used for any selective brain region, as well as for studying internal behaviour of artificial vision systems. In this work, we show results for a novel set of multi-unit recordings from body-selective regions in macaque superior temporal sulcus. Body cells constitute a particularly interesting problem, as different body poses produce vast variability among body images, suggesting a rich set of features driving these neurons. We summarize our contributions as follows:
@@
  • We present results from multi-unit recordings of body-selective neurons in macaque IT cortex, demonstrating that these neurons encode overlapping visual features for bodies and objects,
@@ 2 Related work
A substantial body of work exists on attributing behaviour of deep computer vision models to specific image regions. The most common goal is to understand why classification models output the observed class label, given an input image, meaning that attribution methods are often applied to the very last network layer. Several families of approaches have been proposed for tackling this problem: The first relies on image perturbations like occlusion [12, 13], the second is based on an analysis of latent activations [14, 15], and the third computes gradients of class activations w.r.t. pixel intensities [16, 17, 18]. More recent work [19, 20] has shown that some of these methods rely too strongly on the input image itself, being independent of the network to varying degrees. Even though our reweighting scheme is compatible with any attribution method that is computable for latent activations, we therefore rely on vanilla backpropagation in this work, which has been shown to not be biased towards reconstructing the input image. We provide results for integrated gradients [18] in the appendix. Further, relevant to our work is that of [21] which computes saliency maps for image similarity. The approach is similar to that of Grad-CAM, as it relies on analyzing feature activations before and after a global pooling layer at the end of the network. For that reason, the latter method is mainly suitable for visualizing global similarity as judged by a network trained to do so, rather than more local features relevant for biological neurons along the visual hierarchy in the primate brain.
@@ 3 Methods
3.1 Modelling neural responses
In recent years, a pretrained CNN backbone combined with a trainable linear readout module has become the gold standard in modelling the stimulus-driven variance in neural responses to images [22, 23]. In these models, an image $x\in\mathbb{R}^{3\times h\times w}$ is fed through a CNN $f(\cdot)$ up to a predetermined layer to obtain a latent representation $a\in\mathbb{R}^{c}$. The predicted response for a neuron is then computed as
@@ 3.2 Neuron-specific image similarity
If an image $x_{\text{out}}$ strongly drives a neuron, we propose to interpret the relevant image features by studying a visually similar image from the category that the neuron is selective to. To quantify similarity, we adapt the common method of computing the cosine similarity of latent activations computed using a CNN [24]. Since we have additional information on which latent dimensions drive the neuron, we weight feature $i$ by the corresponding weight $w^{(i)}$. Formally, for images $x_{1}$, $x_{2}$ and their corresponding latent activations $a_{1}:=f(x_{1})$, $a_{2}:=f(x_{2})$, we define the neuron-specific image similarity as
@@ 3.3 Parallel backpropagation for shared-feature visualization
In order to further highlight shared features contributing to model activity, we compute weighted gradients of the latent model activations w.r.t. to image pixels [16]. This method was originally devised for classification models to highlight those image pixels that most strongly influence a given class probability. We build on this work to visualize shared features between images by introducing a simple reweighting procedure. Specifically, we strengthen the influence of features that are shared, while attenuating features that are specific to only one of the images. For the sake of brevity, we illustrate the procedure for the gradients of $x_{\text{out}}$ only, as the process is analogous for $x_{\text{in}}$. First of all, note that the gradient of the predicted neural response $\hat{y}_{\text{out}}$ is given by
@@
Before reweighting the features, we first smooth and normalize each gradient to have unit norm, i.e. $||\frac{\partial}{\partial x_{\text{out}}}a_{\text{out}}^{(i)}||_{2}=1$. Smoothing alleviates pixel-level noise and allows the user to determine regions of pixels with high contribution. It has been shown that a post-processing smoothing step substantially improves the quality of gradient-based attributions [20]. Normalization ensures that the pixel saliency map is not determined solely by the gradient magnitude, as some features may be more sensitive to pixel perturbation than others. More importantly, it allows us to later bound the norm of the saliency map from above, which substantially improves interpretability. Subsequently, we weight each feature by its contribution to the neuron-specific similarity metric $s(\cdot,\cdot)$ given in (2). Formally, we define the weight for feature $i$ as
@@
The revised weight vector $\beta$ imposes high weights only on those features that are highly activated in the feature vectors of both images (captured by $a_{1}^{(i)}$ and $a_{2}^{(i)}$), and are also relevant for the neuron’s activity (captured by $(w^{(i)})^{2}$). Dividing by the product of norms allows us to bound the intensity of the saliency maps by writing
@@ 4 Experimental setup
For gathering neural data, we utilized two sets of images. The first consisted of 475 images of a macaque avatar on a gray background (see Fig. 1, 2). The images were subsampled from a set of 720 images, in which the avatar appeared in 45 unique poses, extracted from nine different behavioural classes, each shown from 16 viewpoints [32]. The second set comprised 6,857 objects from varying categories shown on the same gray background and was combined from the OpenImages dataset [33], as well as several smaller ones [34, 35]. To test for body selectivity, we used an additional set of 2068 control body images including a variety of species. These images were only used to test for category-selectivity of the recorded cells. All other references to body images in the text refer to the original image set showing the monkey avatar. All stimuli were centered with respect to the fixation point. They were shown to the monkey at a resolution of 280x280, and were resized to 224 for the Resnet.
@@
We recorded multi-unit activity (MUA) from and surrounding two fMRI-defined body category-selective patches [5] in the macaque superior temporal sulcus (STS), using 16-channel Plexon V probes, while the subjects performed a fixation task. Since body patch neurons typically discharge sparsely for non-body objects, we employed online stimulus selection: In a first phase, we recorded responses to the set of $475$ monkey avatar images. We then trained a model for each channel to predict neural responses to these images. Subsequently, we predicted neural responses to our set of object images, and for each neuron we selected the highest and lowest predicted activator, as well as the object most similar to the top-activating body image, according to $s(\cdot,\cdot)$. We followed the same procedure for the control body images. Finally, in a second experimental phase we recorded responses to these novel images as well as a subset of 75 of the original body images, to test recording stability. We included recording channels in the analysis if the test/retest stability was higher than .60, as measured by the correlation between responses to body images before and after model fitting. Further, we tested each channel for body-category-selectivity by comparing the median response to the selected objects and the selected control bodies using a Mann-Whitney U-test. Both the animal care and experimental procedures adhere to regional (Flanders) and European guidelines and have been approved by the Animal Ethical Committee of KU Leuven under the protocol number P182/2019. Further experimental details are given in the appendix.
@@ Conclusion.
We presented a method for visualizing selectivity of class-selective visual feature detectors when confronted with out-of-class images. Further, we showed that body-selective neurons encode bodies and objects using an at least partially shared feature set. We visualized these features, providing an explanation for why some objects activate body-selective neurons. Future work could involve using the same method for other category-selective areas, like face patches. Additionally, one could test these visualizations by presenting highlighted fragments to the subject in a closed-loop fashion.
@@ A.2 Additional details for neural data collection
We recorded neuronal responses during passive fixation task using a 2x2 degree fixation window (EyeLink 1000 infrared eye tracker sampling at 1000 Hz). Stimuli were shown gamma-corrected on a 22.5-inch ViewPixx monitor at a distance of about 57 cm, with a resolution of 1920x1080 and a refresh rate of 120 Hz. A set of 475 stimuli of a monkey avatar in various poses were presented 8 times in a pseudorandom sequence for 200 ms on a gray background, with a 250 ms interstimulus interval. During the interstimulus interval, only a fixation dot was present and monkeys could receive a brief juice reward . Stimulus onset and offset were indicated by a photodiode, detecting luminance changes synced with the stimuli, in the display corner (invisible for the monkeys). Control of stimulus presentation, event timing, and juice delivery was managed by an in-house Digital Signal Processing-based computer system, which also monitored the photodiode signal and tracked eye positions. Neuronal data was collected from and surrounding two fMRI-defined body-selective patches in the ventral superior temporal sulcus (STS) using 16-site linear electrodes (Plexon V probe) with Open Ephys acquisition board and software (sampling rate: 30000 Hz, filtered between 500-5000 Hz). Multi-unit activities were extracted using Plexon Offline Sorter, after applying a high-pass Butterworth filter with a cutoff frequency of 250 Hz. From the stimulus event synchronized continuous data, 550 ms trials were extracted, comprising 200 ms prestimulus and 150 ms poststimulus periods. Responsive (at least for one stimulus showing stronger than 5 spikes/s net responses (baseline: -75 to +25 ms, response: +50 to +250 ms)), and body-selective MUAs (p<0.01 Kruskal-Wallis test), with a split-half reliability >0.5 (Spearman Brown corrected) were selected. The responses of these MUA sites were used to predict neural responses to our set of object and body images, and for each neuron we selected the highest and lowest predicted activator, as well as the object most similar to the top-activating avatar stimuli, according to $s(\cdot,\cdot)$. Finally, in a second experimental phase we recorded responses to these object and body images as well as a subset of the original monkey avatar stimuli, to test recording stability (same experimental design as in the first phase). For all neurons considered in this work, we tested for body-selectivity by comparing the median response to bodies to the median response to objects using a Mann-Whitney U-test, considering only channels for which the test detected a significant difference.
@@
Two male 7 years old rhesus monkeys (Macaca mulatta), weighing 9.2 and 11 kg, respectively, contributed to this study. The animals were housed in enclosures at the KU Leuven Medical School and experienced a natural day-night cycle. Each monkey shared its enclosure with at least one other cage companion. On weekdays, dry food was provided ad libitum, and the monkeys obtained water, or other fluids, during experiments until they were satiated. During weekends, the animals received water along with a mixture of fruits and vegetables. The animals had continuous access to toys and other forms of enrichment. After fMRI scanning, we implanted a custom-made plastic recording chamber, allowing a dorsal approach to temporal body patches. In each animal, the location of the recording chamber was guided by the fMRI body localizer described in [40]. Surgery was performed using standard aseptic procedures and under full anesthesia.